# A meshless method for computational electromagnetics with improved dispersion properties

### v1

@@ Abstract
The finite difference time domain method is one of the most popular methods in computational electromagnetics. This work considers two possible ways of generalising it to a meshless setting by employing local radial basis function interpolation. The resulting methods turn out to be unstable, but can be stabilised by adding properly chosen hyperviscosity terms to the update equations. We demonstrate that the proposed meshless methods are convergent and can enjoy a decreased dispersion anisotropy compared to the finite difference time domain method.
@@ 1 Introduction
Accurate simulations of electromagnetic waves are of great interest nowadays with one of the key drivers at the moment being the field of telecommunications, as the carrier frequencies for future communication protocols (e.g. 6G) are expected to be higher, demanding physically accurate simulations that are able to describe increasingly finer geometrical objects, such as antennas. In order to have such a description, we must solve the equations that describe the physics of electromagnetism - Maxwell’s equations. Maxwell’s equations consist of a coupled system of Partial Differential Equations (PDEs) that are too complex to be solved analytically in realistic setups and a numerical procedure must be invoked.
Most of the commonly used methods in computational electromagnetics have been around for decades with several landmark papers appearing in the 1960s. In 1966, for instance, the well known Finite Difference Time Domain (FDTD) method was proposed by Yee [35]. In the same year, the Method of Moments (MoM) was popularised by Richmond [18]. The Finite Element Method (FEM), which can be traced all the way to the work of Courant in 1943 [4], has also been applied to electromagnetics by Silvester in 1969 [24]. A comprehensive review and comparison of these can be found in any of the major textbooks on computational electromagnetics, such as [5]. Among the listed methods, the FDTD is perhaps the simplest to implement and therefore a common starting point when dealing with a given problem. Its simplicity lies in the fact that the FDTD is, in essence, just the application of the usual Finite Difference Method (FDM) on a staggered grid. This approach, however, has its limitations - it is limited to regular, grid-like geometries, making it difficult to efficiently satisfy our requirement of describing fine geometrical objects. Additionally, the underlying grid makes the numerical dispersion relation anisotropic, even if the propagation medium is isotropic.
@@
For that purpose, we will adopt meshless methods, which provide the means of approximation on scattered nodes without any additional structure. Meshless or meshfree methods encompass a variety of different methods, some of which are described in, for example [9]. For our work we opt for methods based on Radial Basis Functions [3] (RBFs). RBFs have first appeared in the context of scattered data interpolation in the 1970s [13] and differential operator approximation in the 1990s [14]. In the 2000s a local RBF-based operator approximation method was introduced [30], which directly generalises the FDM, sparking an interest from the wider scientific community and resulting in a culmination of research in this field, both theoretical and experimental [31, 2].
@@
Having the machinery of meshless methods, a generalisation of the FDTD is then straightforward - we can simply replace the Finite Differences (FD) appearing in the FDTD by their appropriate meshless variant. We will see that there are two natural ways of doing so, resulting in two different generalisations to consider.
@@
At this point, a natural question could arise - why should one bother with a meshless variant of the FDTD, if FEM already solves Maxwell’s equations (albeit in weak form) and is known to be geometrically flexible? Our motivation for considering meshless methods over FEM is twofold. Firstly, unlike meshless methods, FEM requires an underlying triangulation, a generation of which can be tedious and cannot be fully automated, especially in higher dimensions [11]. Secondly, FEM is a fairly complex method - the attractive point of the FDTD is its simplicity and our aim is to preserve this property in the geometrically flexible version. Simplicity of the method can be of great importance when dealing with problems of increasing complexity - not only with regards to the problem geometry but also, for example, inclusion of non-homogenous or even non-linear materials. Furthermore it is also easier to build upon from an implementation perspective, in particular when it comes to parallelisation [5]. Indeed, we will see that, just like the FDTD, our proposed methods will be fully explicit.
@@
It should be mentioned that computational electromagnetics has already been studied in a meshless context. Most of existing work has been focused on weak-form meshless methods [20, 7, 6], although some of the research has been done also using the Method of Fundamental Solution (MFS) [34], which is closely related to the previously mentioned MoM method. As mentioned, however, we are interested in strong-form methods, which resemble FDTD. The latter have also been already investigated in a meshless context [36, 22, 21] by replacing the spatial derivatives appearing in the FDTD by their meshless variants in a similar way as we intend to. However, our proposed method differs from these by two key points, which all of the listed work has in common. Firstly, the listed work staggers the $\vec{E}$ and $\vec{H}$ fields on different nodes, as in the usual FDTD. This staggering can become non-trivial in a scattered setting and would require, for example, a Voronoi tessellation, which some would argue is not truly a meshless method. The second key point is much more important - despite the methods being meshless in their formulations, they seemed to have only been tested on regular node layouts, where the methods can be much better behaved, particularly in terms of stability. In our approach we will abandon the staggering of the $\vec{E}$ and $\vec{H}$ fields and one of our goals is to demonstrate that our methods work also on scattered, irregular nodes.
@@
The paper is structured as follows: In Section 2, the governing equations of electromagnetism are listed. In Section 3, the FDTD is briefly described, along with its previously mentioned shortcomings. In Section 4, RBF interpolation-based meshless methods are introduced. In Section 5, our proposed methods are presented and analysed. In Section 6, the dispersive properties of the proposed methods are also analysed and compared to the FDTD. We then conclude with Section 7.
@@ 2 Governing equations
where $\epsilon_{0}$ and $\mu_{0}$ are the vacuum permittivity and permeability, respectively. Forthcoming analyses will be performed on an effectively 2-dimensional setup, where the fields $\vec{E}$ and $\vec{H}$ only depend on $(x,y)$ coordinates. For this case, it turns out [12] that the above system decouples in two systems, corresponding to the so-called TMz and TEz modes - Transverse Magnetic and Electric, respectively, in the z-direction.
TMz mode considers $\vec{E}(x,y)=(0,0,E_{z}(x,y))$, $\vec{H}=(H_{x}(x,y),H_{y}(x,y),0)$, for which the equations simplify to:
@@ 3 The Finite Difference Time Domain
One of the simplest and most widespread algorithms for solving the aforementioned equations is the Finite Difference Time Domain (FDTD) method, also known as the Yee algorithm. As the name implies, the algorithm discretises the curl equations using the usual Finite Difference (FD) schemes. What sets FDTD algorithm apart from the Finite Difference Method (FDM) is the use of a staggered grid (in this context known as Yee grid) - different components of $\vec{E}$ and $\vec{H}$ fields are discretised at different points in time and space. In the case of a TMz mode the field values are discretised as follows:
@@ 3.1 Shortcomings of FDTD
The FDTD has two main shortcomings that we aim to address with the proposed methods.
The first could already be observed in Figure 1 - a grid-based method has trouble describing irregular boundaries, and the domain has to terminate in a staircase manner. Such a poor boundary description can greatly affect accuracy in more realistic geometries and often, the grid is forced to be made increasingly finer, resulting in very long computational times. Remedies for that, such as the contour FDTD exist [5], however they require manual intervention and can become cumbersome for increasingly complex and realistic domains.
The second shortcoming can be seen on Figure 2, where we simulate the effect of a Gaussian point hard [^2] Hard source at a point means we simply set the field value to a desired source function at each iteration. source, $E_{z}(n)\propto e^{-(n-100)^{2}/2}$, where $n$ is the timestep. We can see numerical dispersion appear in the form of additional, unphysical waves behind the main wave front. The interesting part is the fact that dispersion is angle dependent - with more dispersion in the $x,y$ directions and less on the $y=\pm x$ diagonals.
Figure 2: Snapshot of the FDTD simulation with the described source.
It can be derived [12] that the numerical dispersion relation in FDTD is
@@
where the wave vector is $\vec{k}=(k_{x},k_{y})=k(\cos\phi,\sin\phi)$. As observed, the dispersion relation is anisotropic and depends on the angle of propagation $\phi$, even though the underlying physics is isotropic - the physically correct dispersion relation is well-known: $\omega=ck$. The cause of the unphysical anisotropy is the presence of a grid, which breaks the symmetry in a sense that the $x$ and $y$ directions are special, despite the fact that the curl operator appearing in Maxwell’s equations is rotationally symmetric.
@@ 4 Computation on scattered nodes
To improve the flexibility of the numerical method and ease the description of irregular boundaries, we can abandon the idea of a grid and allow the nodes to be positioned irregularly, such as in Figure 3. To obtain such a set of nodes we use a scattered node generation algorithm, described in [25], which efficiently discretises a given computational domain, producing nodes with the internodal spacing of approximately $h$ (a chosen parameter). The algorithm can also discretise more complex domains, such as CAD models [8]. We will denote the nodes of the resulting discretisation as $\mathbf{x_{i}}=(x_{i},y_{i})$, where $i$ ranges from $1$ to $N$, the discretisation size.
@@ 4.1 Interpolation on scattered nodes
In order to operate on scattered nodes, we need to invoke some tools from meshless numerical analysis. Namely, we will employ Radial Basis Functions (RBFs) to approximate our unknowns, which we now briefly describe. A more in-depth explanation can be found in [9, 16].
We start by introducing the RBF interpolation. We are looking for an approximation of a function $f:\Omega\to\mathbb{R}$ that has known values $f_{i}=f(\mathbf{x}_{i})$ on a scattered node set $X=\{\mathbf{x}_{i}\}_{i=1}^{N}$. The $N$ radial basis functions are defined by our discretisation as: $\Phi_{i}(\mathbf{x})=\varphi(\|\mathbf{x}-\mathbf{x}_{i}\|)$, for some chosen radial basic function $\varphi$. The RBF interpolant is then:
@@
Traditionally, $\varphi$ is chosen to be a strictly positive definite function (as are many popular choices, such as the Gaussians or the Multiquadrics). In that case, the interpolation problem can be shown to admit a unique solution [9]. In practice, to achieve a certain degree of polynomial reproduction we also add monomials to the interpolant:
@@
where $\psi_{i}(x)=(\tilde{A}^{-1}(\Phi(\mathbf{x})\ p(\mathbf{x}))_{i}$, is the $i$-th cardinal function [^4] Also known as Lagrange function. for the interpolation problem and satisfies $\psi_{i}(\mathbf{x}_{j})=\delta_{ij}$.
So far, we have described a global method - the interpolation problem was solved over the whole domain. In practice, the number of nodes $N$ is big and this quickly becomes computationally inefficient. For that reason, we are usually interested in local approximations near a given point $\mathbf{x}_{c}$. In this case we simply write the interpolant (8) over a certain neighbourhood of the point $\mathbf{x}_{c}$, commonly known as its stencil. A stencil is usually taken to simply consist of $n$ ($n\ll N$) nodes that are closest to $\mathbf{x}_{c}$. An example of a stencil can be seen on Figure 3. If we denote the index set of the stencil as $S(\mathbf{x}_{c})$, the local interpolant becomes:
@@
where the unknown coefficients depend on the center point $\mathbf{x}_{c}$ and can be obtained by solving a linear system, as before. It can be shown that for a sufficiently smooth $f$ the error of the interpolation scales as $\mathcal{O}(h^{m+1})$ [31].
Figure 3: An example scattered discretisation, with the nodal spacing of approximately $h$. The domain boundary $\partial\Omega$ and a stencil for a chosen node are also displayed.
@@ 4.2 Differential operator approximation
The second approach is not as widely used in the meshless community, although it is also based on RBF interpolant and has been sucessfully applied to non-trivial problems [33, 28]. It requires us to have a way of approximating $\mathcal{L}$ with the usual finite differences - we have a finite difference stencil given by coefficients $c=(c_{1},\dots,c_{k})$ and offsets $o=(\mathbf{o}_{1},\dots,\mathbf{o}_{k})$. For example, for the central derivative in the $x$ direction we have c=(0,1,1)c=(0,-1,1) and o=((0,0),(1,0),(1,0))o=((0,0),(-1,0),(1,0)), while for the 5-point Laplace operator approximation we have c=(4,1,1,1,1)c=(-4,1,1,1,1), o=((0,0),(1,0),(1,0),(0,1),(0,1))o=((0,0),(1,0),(-1,0),(0,1),(0,-1)). Approximating the operator $\mathcal{L}$ in the finite difference manner then amounts to:
@@
Figure 4: A visual demonstration of the virtual stencil approach for the case of a 5-point stencil with a spacing $\delta$.
where $l$ is the order of the operator $\mathcal{L}$ and $\delta$ the spacing between the nodes. While this works well on a grid, the problem in a scattered layout is that $\mathbf{x}_{c}+\delta\mathbf{o}_{i}$ is in general not in our discretisation, so we do not have access to the value of $f$ at this point. We circumvent this by using the RBF interpolant, effectively imagining a virtual stencil around the point $x_{c}$, as seen on Figure 4:
@@
where $w^{\mathcal{L}}_{j}=\delta^{-l}\sum_{i=1}^{k}c_{i}\psi_{j}^{c}(\mathbf{x}_{c}+\delta\mathbf{o}_{i})$, are the differentiation weights, expressing the approximation of the differential operator $\mathcal{L}$ and can again be understood as the generalised finite difference coefficients, with the offsets given by the stencil nodes. The virtual stencil spacing $\delta$ is a parameter of the method an usually expressed in the units of the internodal spacing $h$ as $\delta=\sigma h$. Experiments show that it is desirable to have $\sigma<1$ [19, 33]. As far as the authors know, the second approach does not have a well-established name and we will refer to it as Radial Basis Function-generated Virtual Finite Differences (RBF-VFD) in this paper. The order of the RBF-VFD method can be easily deduced: The interpolation step has order $\mathcal{O}(h^{m+1})$ and the interpolant is divided by $h^{l}$, bringing the order to $\mathcal{O}(h^{m+1-l})$. The second step, the application of the chosen finite difference stencil has a fixed, known order $\mathbf{ord}$. The total order is then simply $\mathrm{min}(m+1-l,\mathbf{ord})$.

### v2

# An RBF-based method for computational electromagnetics with reduced numerical dispersion

@@ Abstract
The finite difference time domain method is one of the simplest and most popular methods in computational electromagnetics. This work considers two possible ways of generalising it to a meshless setting by employing local radial basis function interpolation. The resulting methods remain fully explicit and are convergent if properly chosen hyperviscosity terms are added to the update equations. We demonstrate that increasing the stencil size of the approximation has a desirable effect on numerical dispersion. Furthermore, our proposed methods can exhibit a decreased dispersion anisotropy compared to the finite difference time domain method.
keywords:
meshless , electromagnetics , FDTD , RBF , hyperviscosity , dispersion
@@ 1 Introduction
Accurate simulations of electromagnetic waves have been of great interest lately with one of the key drivers being the field of telecommunications, as the carrier frequencies for future communication protocols (e.g. 6G) are expected to be higher, demanding physically accurate simulations that are able to describe increasingly finer geometrical objects, such as antennas. In order to satisfy that requirement, we must solve the equations that describe the physics of electromagnetism – Maxwell’s equations. Maxwell’s equations form a coupled system of first order Partial Differential Equations (PDEs) that are too complex to be solved analytically in realistic setups. In order to obtain the solution a numerical procedure must be invoked.
Most of the commonly used methods in computational electromagnetics have been around for decades with several landmark papers appearing in the 1960s. In 1966, for instance, the well known Finite Difference Time Domain (FDTD) method was proposed by Yee [37]. In the same year, the Method of Moments (MoM) was popularised by Richmond [18]. The Finite Element Method (FEM), which can be traced all the way to the work of Courant in 1943 [4], has also been applied to electromagnetics by Silvester in 1969 [25]. A comprehensive review and comparison of these can be found in any of the major textbooks on computational electromagnetics, such as [5]. Among the listed methods, the FDTD is arguably the simplest to implement and therefore a common starting point when dealing with a given problem. Its simplicity lies in the fact that the FDTD is, in essence, just the application of the usual Finite Difference Method (FDM) on a staggered grid. This approach, however, has its limitations – it is limited to regular, grid-like geometries, making it difficult to efficiently satisfy our requirement of describing fine geometrical objects. Additionally, the underlying grid makes the numerical dispersion relation anisotropic, even if the propagation medium is isotropic.
@@
For that purpose, we will adopt meshless methods, which provide the means of function approximation on scattered nodes without any additional structure. The terms “meshless” or “meshfree” encompass a variety of different methods, some of which are described in, for example [8]. For our work we opt for methods based on Radial Basis Functions (RBFs) [3]. RBFs first appeared in the context of scattered data interpolation in the 1970s [13] and were applied to differential operator approximation in the 1990s [14]. In the 2000s a local RBF-based operator approximation method was introduced [31], which directly generalises the Finite Difference Method (FDM), sparking interest from the wider scientific community and resulting in a culmination of research in the field, both theoretical and experimental [32, 2].
@@
With the machinery of meshless methods on hand, a generalisation of the FDTD is then straightforward – we can simply replace the Finite Differences (FD) appearing in the FDTD by their appropriate meshless variants. We will see that there are two natural ways of doing so, resulting in two different generalisations to consider.
@@
At this point a natural question may arise - why should one bother with a meshless variant of the FDTD when FEM already solves Maxwell’s equations (albeit in weak form) and is known to be geometrically flexible? Our motivation for considering meshless methods over FEM is twofold. Firstly, unlike meshless methods, FEM requires an underlying mesh, a generation of which can be tedious and cannot be fully automated, especially in higher dimensions [11]. Secondly, FEM is a fairly complex method – the attractive point of the FDTD is its simplicity and our aim is to preserve this property in the geometrically flexible version. Simplicity of the method can be of great importance when dealing with problems of increasing complexity – not only with regards to the problem geometry but also, for example, inclusion of non-homogenous or even non-linear materials. Another aspect of great practical value is parallelisation, which is easy to implement in an explicit method, such as the FDTD, and generally difficult in FEM [5].
@@
It should be mentioned that computational electromagnetics has already been studied in meshless context. Much of existing work has been focused on weak-form meshless methods [21, 7, 6], although some research has been done also using the Method of Fundamental Solution (MFS) [36], which is closely related to the previously mentioned method of moments. However, our interests lie in strong-form methods, just like the FDTD. Existing work has been done also in this case [38, 23, 22], where different authors adopted a similar idea of replacing the spatial derivatives appearing in the FDTD by their meshless variants in a similar way as we intend to. However, our proposed methods differ from all of the listed work by two key points. Firstly, previous authors have always staggered the $\vec{E}$ and $\vec{H}$ fields on different nodes, as in the usual FDTD. This staggering can become non-trivial in a scattered setting and would require, for example, a Voronoi tessellation, which some would argue is not truly a meshless method. The second key point is perhaps more important – despite the methods being meshless in their formulations, they seemed to have only been tested on regular node layouts, where the methods can be much better behaved, particularly in terms of stability. In our approach we abandon the staggering of the $\vec{E}$ and $\vec{H}$ fields and our main goal is to demonstrate that our methods work also on scattered, irregular nodes.
@@
The paper is structured as follows: In Section 2, the governing equations of electromagnetism are listed. In Section 3, the FDTD is briefly described, along with its previously mentioned shortcomings. In Section 4, RBF interpolation-based meshless methods are introduced. In Section 5, our proposed methods are introduced. In Section 6, we analyse the numerical dispersion of our proposed methods and compare it to the FDTD. We then conclude with Section 7.
@@ 2 Governing equations
where $\epsilon_{0}$ and $\mu_{0}$ are the vacuum permittivity and permeability, respectively. Forthcoming analyses will be performed on an effectively 2-dimensional setup [^1] The proposed methods admit a straightforward generalisation to 3D., where the fields $\vec{E}$ and $\vec{H}$ only depend on $(x,y)$ coordinates. For this case, it turns out [12] that the above system decouples in two systems, corresponding to the so-called TMz and TEz modes – transverse magnetic and electric, respectively, in the $z$ direction.
TMz mode considers $\vec{E}(x,y)=(0,0,E_{z}(x,y))$ and $\vec{H}=(H_{x}(x,y),H_{y}(x,y),0)$, for which the equations simplify to:
@@ 3 The Finite Difference Time Domain
One of the simplest and most widespread algorithms for solving the aforementioned equations is the Finite Difference Time Domain (FDTD) method, also known as the Yee algorithm. As the name implies, the algorithm discretises the curl equations using the usual Finite Difference (FD) schemes. What sets FDTD algorithm apart from the Finite Difference Method (FDM) is the use of a staggered grid (in this context known as a Yee grid) – different components of $\vec{E}$ and $\vec{H}$ fields are discretised at different points in time and space. In the case of a TMz mode the field values are discretised as follows:
@@ 3.1 Shortcomings of FDTD
The FDTD has two main shortcomings that we aim to address with the proposed methods. The first could already be observed in Figure 1 – a grid-based method has trouble describing irregular boundaries, and the domain has to terminate in a staircase manner. Such a poor boundary description can greatly affect accuracy in more realistic geometries and often, the grid is forced to be made increasingly finer, resulting in very long computational times. Remedies for that, such as the contour FDTD exist [5], but they require manual intervention and can become cumbersome for increasingly complex domains.
The second shortcoming can be seen on Figure 2, where we simulate the effect of a hard [^3] Hard source refers to setting the field values to a desired source function at each timestep. Gaussian point source, $E_{z}(n)\propto e^{-(n-100)^{2}/2}$, where $n$ is the timestep. We can see numerical dispersion appear in the form of additional, unphysical waves behind the main wave front. The interesting part is the fact that this dispersion is angle dependent – with more dispersion in the $x,y$ directions and less on the $y=\pm x$ diagonals.
Figure 2: Snapshot of the FDTD simulation at $n=400$, demonstrating anisotropic dispersion relation.
In fact, FDTD numerical dispersion relation can be derived [12] and equals
@@
where the wave vector is $\vec{k}=(k_{x},k_{y})=k(\cos\phi,\sin\phi)$. As observed, the dispersion relation is anisotropic and depends on the angle of propagation $\phi$, even though the underlying physics is isotropic – a physically correct dispersion relation is well-known: $\omega=c\|\vec{k}\|$. The cause of this unphysical anisotropy is the presence of a grid, which breaks the symmetry in a sense that the $x$ and $y$ directions are special, despite the fact that the curl operator appearing in Maxwell’s equations is rotationally symmetric.
@@ 4 Computation on scattered nodes
To improve the flexibility of the numerical method and ease the description of irregular boundaries, we abandon the idea of a grid and allow the nodes to be positioned irregularly, such as in Figure 3. To obtain such a set of nodes we use a scattered node generation algorithm, described in [26], which efficiently discretises a given computational domain, producing nodes with the internodal spacing of approximately $h$ (a chosen parameter). We will denote the nodes of the resulting discretisation as $\mathbf{x_{i}}=(x_{i},y_{i})$, where $i$ ranges from $1$ to $N$, the discretisation size.
@@ 4.1 Interpolation on scattered nodes
In order to operate on scattered nodes, we need to invoke some tools from meshless numerical analysis. Namely, we will employ Radial Basis Functions (RBFs) to approximate our unknowns. What follows is a brief description of the methods employed. A more in-depth explanation can be found in [8, 16].
We start by introducing RBF interpolation. We are looking for an approximation of a function $f:\Omega\to\mathbb{R}$ that has known values $f_{i}=f(\mathbf{x}_{i})$ on a scattered node set $X=\{\mathbf{x}_{i}\}_{i=1}^{N}$. The $N$ radial basis functions are defined by our discretisation as: $\Phi_{i}(\mathbf{x})=\varphi(\|\mathbf{x}-\mathbf{x}_{i}\|)$ for some chosen radial basic function $\varphi$. The RBF interpolant is then:
@@
Usually, $\varphi$ is chosen to be a strictly positive definite function (as are many popular choices, such as Gaussians or multiquadrics). In this case, the interpolation problem can be shown to admit a unique solution [8]. In practice, to achieve a certain degree of polynomial reproduction we augment our interpolant with monomials:
@@
where $\psi_{i}(x)=(\tilde{A}^{-1}(\Phi(\mathbf{x})\ p(\mathbf{x}))_{i}$, is the $i$-th cardinal function [^5] Also known as Lagrange function. and satisfies $\psi_{i}(\mathbf{x}_{j})=\delta_{ij}$.
So far, we have described a global method – the interpolation problem was solved over the whole domain. In practice, number of nodes $N$ is large and this quickly becomes computationally inefficient and ill-conditioned. For that reason, we are usually interested in local approximations near a given point $\mathbf{x}_{c}$ and we write the interpolant (8) over a certain neighbourhood of the point $\mathbf{x}_{c}$, commonly known as its stencil. A stencil is usually taken to simply consist of $n$ ($n\ll N$) nodes that are closest to $\mathbf{x}_{c}$ – an example can be seen on Figure 3. If we denote the index set of the stencil as $S(\mathbf{x}_{c})$, the local interpolant becomes:
@@
where the unknown coefficients depend on the center point $\mathbf{x}_{c}$ and can be obtained by solving a linear system, as before. It can be shown that for a sufficiently smooth $f$ the pointwise interpolation error scales as $\mathcal{O}(h^{m+1})$ [32].
Figure 3: An example scattered discretisation, with the nodal spacing of approximately $h$. The domain boundary $\partial\Omega$ and the stencil for a chosen node are also displayed.
@@ 4.2 Differential operator approximation
The second approach is not as widely used in the meshless community, although it is also based on RBF interpolantion and has been sucessfully applied to non-trivial problems [35, 29]. It requires us to have a way of approximating $\mathcal{L}$ with the usual finite differences – we have a finite difference stencil given by coefficients $c=(c_{1},\dots,c_{k})$ and offsets $o=(\mathbf{o}_{1},\dots,\mathbf{o}_{k})$. For example, for the central derivative in the $x$ direction we have c=(0,1,1)c=(0,-1,1) and o=((0,0),(1,0),(1,0))o=((0,0),(-1,0),(1,0)), while for the 5-point Laplace operator approximation we have c=(4,1,1,1,1)c=(-4,1,1,1,1), o=((0,0),(1,0),(1,0),(0,1),(0,1))o=((0,0),(1,0),(-1,0),(0,1),(0,-1)). Approximating the operator $\mathcal{L}$ in the finite difference manner then amounts to:
@@
Figure 4: A visual demonstration of the virtual stencil approach for the case of a 5-point stencil with spacing $\delta$.
where $l$ is the order of the operator $\mathcal{L}$ and $\delta$ the spacing between the nodes. While this works well on a grid, the problem is that in a scattered layout $\mathbf{x}_{c}+\delta\mathbf{o}_{i}$ is in general not in our discretisation, so we do not have access to the value of $f$ at this point. We circumvent this by using the RBF interpolant instead, effectively imagining a virtual stencil around the point $x_{c}$, as seen on Figure 4:
@@
where $w^{\mathcal{L}}_{j}=\delta^{-l}\sum_{i=1}^{k}c_{i}\psi_{j}^{c}(\mathbf{x}_{c}+\delta\mathbf{o}_{i})$, are the differentiation weights expressing the approximation of the differential operator $\mathcal{L}$ and can again be understood as the generalised finite difference coefficients. The virtual stencil spacing $\delta$ is a parameter of the method and usually expressed in the units of the internodal spacing $h$ as $\delta=\sigma h$. Experiments show that it is desirable to have $\sigma<1$ [19, 35]. As far as the authors know, this second approach does not have a well-established name and we will refer to it as Radial Basis Function-generated Virtual Finite Differences (RBF-VFD) in this paper. The order of the RBF-VFD approximant can be easily deduced: error of the interpolation step scales as $\mathcal{O}(h^{m+1})$ and the interpolant is divided by $h^{l}$, reducing the exponent to $\mathcal{O}(h^{m+1-l})$. The second step, the application of the chosen finite difference stencil has a fixed, known order $\mathbf{ord}$. The total order is then simply $\min(m+1-l,\mathbf{ord})$.